OpenStudy (anonymous):

Solve the system of equations by graphing. 2x-y=2 , 5x+y=5

OpenStudy (jdoe0001):

so graph it :)

OpenStudy (reemii):

you have to draw two lines and look at the intresection. did you do something already?

OpenStudy (anonymous):

i just don't know the two plotting points for each equation :/

OpenStudy (reemii):

take one equation and make it look like "\(y = ax + b\)". For example the first equation becomes \(y=2x-2\). This should enable you to draw the line.

OpenStudy (anonymous):

would it be (0,-2) (0,2)?

OpenStudy (anonymous):

for the first one

OpenStudy (reemii):

oops. "y=2x-2" so if x=1, then y=2-2=0.

OpenStudy (anonymous):

so then what would be the two plotting points for the first one?

OpenStudy (anonymous):

@reemii

OpenStudy (reemii):

to draw the line you have to methods 1) find two points of the line and draw it 2) "y=2x-2" tells that the slope is equal to 2 and the line cuts the y-axis at (0,-2). you do the first method. when x=0 you get y=-2, when x=1, y=0. so the points are (0,-2) and (1,0).

OpenStudy (reemii):

there's no prefered choice for taking the two points, you just have to choose 2 different values for \(x\) and find the corresponding \(y\). It is often simple to take \(x=0\) and \(x=1\).

OpenStudy (reemii):

Is it working well for the other equation?

OpenStudy (anonymous):

the second one would be y=5x+5?

OpenStudy (anonymous):

@reemii

OpenStudy (reemii):

there's a minus sign missing.

OpenStudy (anonymous):

so then it would be y=5x-5?

OpenStudy (reemii):

rather \(y=-5x+5\). The operation to obtain it was to -remove 5x- on both sides. Now you get two points and you'll be able to draw the second line.

OpenStudy (anonymous):

oh ok. so the plotting points for this one would be (0,5) (0,-5)

OpenStudy (reemii):

how do you obtain these points?

OpenStudy (reemii):

explain the process, because one point is ok, the other isn't.

OpenStudy (anonymous):

or would it be (1,0)?

OpenStudy (reemii):

(1,0) yes! and (0,-5) was wrong.

OpenStudy (anonymous):

:) so (1,0) and (0,5)?

OpenStudy (anonymous):

@reemii

OpenStudy (reemii):

that's right. can you finish the problem?

OpenStudy (anonymous):

Thank you so much! yes i did :)

OpenStudy (reemii):

congrats ;)